### v1

# On the identifiability of causal graphs with multiple environments

Abstract
Causal discovery from i.i.d. observational data is known to be generally ill-posed. We demonstrate that if we have access to the distribution of a structural causal model, and additional data from only two environments that sufficiently differ in the noise statistics, the unique causal graph is identifiable. Notably, this is the first result in the literature that guarantees the entire causal graph recovery with a constant number of environments and arbitrary nonlinear mechanisms. Our only constraint is the Gaussianity of the noise terms; however, we propose potential ways to relax this requirement. Of interest on its own, we expand on the well-known duality between independent component analysis (ICA) and causal discovery; recent advancements have shown that nonlinear ICA can be solved from multiple environments, at least as many as the number of sources: we show that the same can be achieved for causal discovery while having access to much less auxiliary information.
1 Introduction
Causal discovery seeks to recover cause–effect structure from data, which allows counterfactual reasoning and prediction under interventions (Pearl, 2009; Peters et al., 2017; Spirtes, 2010; Spirtes et al., 2000). However, learning causal structure from purely observational i.i.d. data is, in general, ill-posed: multiple directed acyclic graphs (DAGs) are distributionally equivalent, i.e., indistinguishable from the data distribution.
In the interventional causal discovery literature, hard interventions—directly modifying the causal structure—are known to unlock identifiability of the underlying graph. Classic results from Eberhardt et al. (2005) show that the number of sufficient hard interventions to identify the causal order scales logarithmically with the number of nodes. A large corpus of intervention-based causal discovery research has largely built on these findings (Eberhardt, 2008; He & Geng, 2008; Hauser & Bühlmann, 2012; Shanmugam et al., 2015; Kocaoglu et al., 2017; Wang et al., 2017; Lindgren et al., 2018; Eaton & Murphy, 2007; Triantafillou & Tsamardinos, 2015; Lorch et al., 2022; Ke et al., 2023b).
Recent work has explored the problem of causal graph identifiability from multiple environments and soft interventions (i.e., in the setting where non i.i.d. data might naturally occur and does not stem from changes in the causal structure) (Perry et al., 2022; Huang et al., 2020; Heinze-Deml et al., 2018; Peters et al., 2015; Ghassami et al., 2017, 2018; Jaber et al., 2020; Jalaldoust et al., 2025; Brouillard et al., 2020; Heurtebise et al., 2025): however, from an identifiability perspective, these results do not provide guarantees of recovery of the unique causal graph with a limited number of environments under generic assumptions.
Our research overcomes this limitation. We prove that, for structural causal models (SCMs) with arbitrary nonlinear mechanisms, auxiliary information from only two sufficiently distinct environments is enough to identify the unique causal graph. Our only constraint is the Gaussianity of the noise terms; however, we outline potential ways to relax this requirement. To our knowledge, this is the first proof of identifiability for full graphs of arbitrary size and generic functional mechanisms from a constant number of environments. Strengthening our findings is the contrast with hard-intervention regimes, where the number of experiments needs to scale with the number of nodes.
Our work is also of independent methodological interest. In particular, key to our theory is the duality between causal discovery and independent component analysis (ICA). Reizinger et al. (2023) recently formalized that nonlinear ICA identifiability results naturally extend to structure learning (well known in the linear case since Shimizu et al. (2006)). This is of great relevance in light of the late advancements in multi-environment ICA identifiability pioneered by Hyvärinen & Morioka (2016); however, directly bootstrapping these findings to causal discovery doesn’t carry great promise, being ICA the harder problem of the two: we show that where ICA identifiability requires a number of environments that scales linearly with the number of variables, causal graph identifiability can be achieved with data from just two extra domains. This calls for causality-only identifiability results in the multi-environment setting, as developed in our work. Inspired by the recent success of ICA with multiple environments, we are hopeful that our approach paves the way to novel causality theory that weakens the requirements in terms of heterogeneity of the data and parametric assumptions.
Our main contributions are summarized as follows:
  • We show that the causal graph underlying an arbitrary invertible causal model with Gaussian noise terms is identifiable from only two auxiliary environments, when they sufficiently vary. Moreover, we outline potential avenues to relax the Gaussianity assumption.
  • A methodological contribution consisting of proof techniques that are novel for causal discovery and leverage the (well-known) duality between structural causal models and independent component analysis; to the best of our knowledge, these are the first causality-only identifiability results for nonlinear SCMs that stem from this connection.
  • We empirically validate our theory. Our synthetic experiments on bivariate models reflect that when the assumptions of our theory are met we can infer the causal direction, even for cases that were previously known to be non-identifiable.
2 Related works
Soft interventions and multiple environments for causal discovery.
Several works in the literature have addressed causal discovery identifiability and estimation via non i.i.d. data (stemming from soft interventions and multiple environments). Peters et al. (2015) and Heinze-Deml et al. (2018) identify the parents of a designated target node via invariance across environments, yielding partial identifiability of causal directions. Huang et al. (2020) use nonstationarity to recover the skeleton and orient some edges. Perry et al. (2022) leverage sparse mechanism shifts, proving high-probability graph recovery with bounds that improve as the number of environments grows. Ghassami et al. (2017, 2018) and Heurtebise et al. (2025), similarly to our work, study identifiability of structural causal models from multiple environments, but their identifiability results are specialized to the linear case. Recently, Jalaldoust et al. (2025) formulated a statistical test that can find a superset of the parents of a target node. Yang et al. (2018); Brouillard et al. (2020); Jaber et al. (2020) characterize equivalence classes identifiability from interventions. From a methodological perspective, Brouillard et al. (2020); Ke et al. (2023a) introduce differentiable approaches to causal discovery with interventions; Mooij et al. (2020) propose a unifying framework for causal discovery from observational and multi-environment data. All of these results are complementary to our work, which is, to the best of our knowledge, the first to provide guarantees of identifiability of the causal graph from a finite number of auxiliary additional environments, potentially only two.
ICA and causal discovery.
The seminal work of Shimizu et al. (2006) shows that if an SCM can be expressed as a linear non-Gaussian ICA model, the underlying causal graph is identifiable. Reizinger et al. (2023) generalize this to the nonlinear case. Monti et al. (2020) show that time contrastive ICA (Hyvärinen & Morioka, 2016) can identify bivariate causal graphs with arbitrary nonlinear mechanism. The common ground of these findings is that they adapt the existing ICA identifiability theory to the problem of causal discovery. This approach is clearly important, especially in the light of the recent advancement in multi-environment ICA identifiability (Hyvärinen et al., 2019; Khemakhem et al., 2020a, b; Gresele et al., 2019; Hälvä & Hyvärinen, 2020; Hyvärinen & Morioka, 2017; Hälvä et al., 2021); however, in the nonlinear setting, it fails to capture the gap between the two problems: while ICA attempts to recover the mixing function and the independent sources at each point, causal discovery concerns the much simpler problem of structure identifiability. Our work shows that this difference is key to demonstrating causal discovery identifiability from a constant number of sufficiently different environments, where ICA requires at least as many as the number of sources (see e.g. Theorem 1 in Hyvärinen & Morioka (2016)).
3 Preliminaries
First, we define structural causal models, independent component analysis, and how they relate. Then, we describe the problem of causal discovery from multiple environments and define identifiability of causal graphs in this context.
3.1 Structural causal models and ICA
Let us consider a set of causal variables $\mathbf{X}$, with components generated according to a structural causal model
$X_{i}:=F_{i}(\mathbf{X}_{\operatorname{PA}_{i}},S_{i}),\hskip 5.0pt\forall i=1,...,d,$ (1)
where $\mathbf{X}_{\operatorname{PA}_{i}}$ are the causes of $X_{i}$, specified by a directed acyclic graph (DAG) $\mathcal{G}$ with nodes $\mathbf{X}$. $\operatorname{PA}_{i}\subset\{1,...,d\}$ denotes the indices of the parents of $X_{i}$ in the graph (see Section˜D.1 for precise definitions on graphs). The functions $F_{i}$ are the causal mechanisms that map causes to effects. We assume mutually independent noise terms $\mathbf{S}=(S_{1},...,S_{d})$ with density $p_{\theta}$, where $\theta$ is a set of parameters defining the density function. Further, we restrict to structural causal models where there are no latent common causes.
Notational remarks.
We use uppercase letters for random variables (or vectors), lowercase letters for their realizations. Vectors are denoted in bold, so that we have $\mathbf{v}=(v_{i})_{i=1}^{d}$, where $v_{i}$ are the scalar vector’s components. Probability density functions are differentiated by their argument, where the distinction is clear from the context: for example, for a random vector $\mathbf{Z}$ we only write $p(\mathbf{z})$ to specify its density at a certain value $\mathbf{z}$. Further, we define the support to keep track of the nonzero entries in matrices: for $M\in\mathbb{R}^{m\times n}$, $\operatorname{supp}(M):=\{(i,j)|i\in[m],j\in[n]\textnormal{ and }M_{ij}\neq 0\}$; for a matrix valued function $M:\mathbb{R}^{d}\to\mathbb{R}^{m\times n}$ the support is defined as $\operatorname{supp}(M)=\{(i,j)|i\in[m],j\in[n]\textnormal{ and there is }\mathbf{x}\in\mathbb{R}^{d}\textrm{ s.t. }M_{ij}(\mathbf{x})\neq 0\}$. Also, we use $d]:=\{1,...,d\}$. Given a vector $\mathbf{V}=(V_{i})_{i\in[d]}$ and a subset $I\subset[d]$, we define $\mathbf{V}_{I}:=(V_{i})_{i\in I}$.
It is well known that the SCM of Equation˜1 can be expressed in the form of an ICA model
$\mathbf{X}=\mathbf{f}(\mathbf{S}),$ (2)
where $\mathbf{f}$ is the ICA mixing function, uniquely specified by the SCM (we show how to construct $\mathbf{f}$ in Section˜D.2).
Definition 1 (ICA model).
We define a pair $(\mathbf{f},p_{\theta})$ as an ICA model, where $\mathbf{f}$ is a diffeomorphism in $\mathbb{R}^{d}$, and $p_{\theta}$ is a factorized density parameterized by $\theta$.
A more detailed introduction to independent component analysis is presented in Appendix˜B.
It is known (Reizinger et al., 2023) that, under some faithfulness assumption, the support of the Jacobian of the mixing function completely identifies the causal structure.
Definition 2 (Faithfulness).
Consider $\mathbf{x}=\mathbf{f}(\mathbf{s})$. We say that $J_{\mathbf{f}^{-1}}(\mathbf{x})$ is faithful if for each $i,j\in[d]$ $J_{\mathbf{f}^{-1}}(\mathbf{x})_{ij}=0\iff$ $S_{i}$ is constant in $X_{j}$ on the entire domain. In other words:
$\operatorname{supp}(J_{\mathbf{f}^{-1}}(\mathbf{x}))=\operatorname{supp}(J_{\mathbf{f}^{-1}}).$ (3)
Proposition 1 (Proposition 1 in Reizinger et al. (2023)).
Let $J_{\mathbf{f}^{-1}}(\mathbf{x})$ faithful. Then, for each $i\neq j$:
$J_{\mathbf{f}^{-1}}(\mathbf{x})_{ij}=0\iff j\not\in\operatorname{PA}_{i}.$
This formulation of faithfulness is well known and at the core of the LiNGAM algorithm for linear SCMs (Shimizu et al., 2006), and is satisfied almost everywhere under some regularity conditions on $\mathbf{f}$. When this is the case, the above proposition means that for causal discovery we are interested in the support of the inverse Jacobian, and, by Equation˜3, this can be recovered by having access to the support at a single point where faithfulness is satisfied.
Next, we introduce the notion of environment and define the causal discovery problem when multiple environments are available.
3.2 Definition of identifiability from multiple environments
Intuitively, causal discovery is the inference problem of finding the causal graph underlying a structural causal model from the data. We are interested in causal discovery from multiple environments, i.e., when data are collected from different but related structural causal models (which we express as ICA models).
Definition 3 (Environment).
Let $\mathbf{\mathbf{X}}=\mathbf{f}(\mathbf{S})$ be an ICA model. Consider the random variable $\mathbf{S}^{i}\sim p^{i}$. For i=1,,ki=1,...,k, we call the ICA model $\mathbf{X}^{i}=\mathbf{f}(\mathbf{S}^{i})$ an auxiliary environment. We adopt the convention that $\mathbf{X}^{0},\mathbf{S}^{0}:=\mathbf{X},\mathbf{S}$, and call $i=0$ the base environment. $p^{i}$ denotes the probability density of the sources defined by the $i^{\textnormal{th}}$ environment.
The key part of our definition is that the mixing function is invariant across environments, while we allow for changes in the sources distribution: if $\mathbf{f}$ is obtained from a structural causal model (as it is assumed across all of our paper), all auxiliary environments share the same causal mechanisms and causal graph as the base model $\mathbf{X}^{0}=\mathbf{f}(\mathbf{S}^{0})$.
Next, we formalize what we mean by identifiability in the context of causal discovery with multiple environments. Intuitively, identifiability is achieved when the graph underlying the structural causal model is uniquely specified by the causal variables’ distribution. In the definition, we denote the pushforward of a density $p$ by $\mathbf{f}$ with $\mathbf{f}_{*}p$.
Definition 4 (Identifiability of the causal graph).
Consider the auxiliary environments $(\mathbf{f},p^{i}_{\theta})$ obtained from the base causal model of Equation˜1, i=1,,ki=1,...,k. Let $\mathcal{F}$ be the space of diffeomorphisms in $\mathbb{R}^{d}$ and $\mathcal{P}$ a family of factorized densities. We say that the causal graph underlying the SCM is identifiable if, given $(\widehat{f},p^{i}_{\widehat{\theta}})\in\mathcal{F}\times\mathcal{P}$, i=1,,ki=1,...,k, then:
$\mathbf{f}_{*}p_{\theta}^{i}=\widehat{\mathbf{f}}_{*}p_{\widehat{\theta}}^{i}\quad\forall i\in[k]\implies\operatorname{supp}(J_{\mathbf{f}^{-1}})=\operatorname{supp}(J_{\widehat{\mathbf{f}}^{-1}}).$
The above definition of identifiability, based on the support of the Jacobian inverse of the mixing function, may be a bit unfamiliar, but it’s equivalent to what is commonly meant when asking that a causal DAG is identifiable: any alternative causal model that matches the distribution of the data is compatible only with the ground truth causal graph (represented with the inverse Jacobian’s support).
Relation with ICA identifiability.
Compare Definition˜4 of identifiability of the causal graph with the notion of identifiability in ICA of Definition˜5 in the appendix: for causal discovery, all we care about is the support of $J_{\mathbf{f}^{-1}}$, which can be identified from any point where the Jacobian is faithful; for independent component analysis, we need to guarantee that the exact values of the Jacobian can be recovered over each point of the domain, up to trivial indeterminacies. This phrasing clarifies that, in the nonlinear setting (where the Jacobian varies with $\mathbf{x}$), causal discovery is a much simpler problem than ICA: it only requires identifying the support at a single point, rather than the value at any point. This is reflected in our main identifiability result (Theorem˜1): we will show that the causal graph of a nonlinear SCM can be identified with the information from only two auxiliary environments; this in stark contrast with ICA identifiability results for general mixing functions, that usually require a number of environments that scales linearly ($\mathcal{O}(d)$) with the number of sources.
Problem definition. We aim to characterize the conditions under which the causal graph $\mathcal{G}$ is identifiable from the fewest possible environments.
4 Theory
To develop our theory, we rely on the following assumptions on the ICA model of Equation˜2.
Assumption 1.
$\mathbf{f}$ is invertible and twice differentiable.
Assumption 2.
Each environment is obtained as a rescaling of $\mathbf{S}$, namely $\mathbf{S}^{i}=L_{i}\mathbf{S}$ for each $i\in[e]$, with $L_{i}=\mathrm{diag}(\lambda_{1}^{i},\dots,\lambda_{d}^{i})$ and $\lambda_{j}^{i}\neq 0$.
Assumption 3.
For $\mathbf{f}^{-1}(\mathbf{x})=\mathbf{s}$ where $\mathbf{s}=\mu_{\mathbf{S}}$, the mean of the vector of sources, the Jacobian is faithful (Definition˜2).
Assumption 4.
$\mathbf{S}$ has Gaussian density $p_{\theta}$ with $\theta$ mean and covariance matrix parameters.
Discussion on the Assumptions 1-4.
˜1 is standard when proving identifiability: the results in Hoyer et al. (2008); Zhang & Hyvärinen (2009); Immer et al. (2022) are based on higher-order derivatives, and have strong requirements that guarantee invertibility. Also ˜2 is mild and somewhat necessary: it simply asks that the interventions are meaningful, i.e. that they affect the variance; interventions on the mean, intuitively, are not informative as they shift the density graph by a constant, without affecting its shape (the gradient and the Hessian of the density, where information about the causal graph lies). ˜3 requires that the Jacobian of the inverse of the mixing function is informative about the causal structure at the mean of $\mathbf{S}$ (and it’s almost surely verified over $\mathbf{X}$ samples, under some generic regularity conditions on $\mathbf{f}$). The reason behind it is that we probe the identifiability of the Jacobian’s support at the mean. The only real simplifying constraint is ˜4 of the Gaussianity of the sources, which is, however, not new in the literature (see e.g.Rolland et al. (2022)). Later, we discuss why this assumption is needed in the paper and potential ways to relax it.
In the remainder of the paper we demonstrate that, under these assumptions, leveraging the ICA formalism we can prove the identifiability of causal graphs, potentially with as few as two auxiliary environments. Our starting point is the invertibility $\mathbf{f}$, so that we can write the density of $\mathbf{X}$ with the change of variable for each value $\mathbf{x}=\mathbf{f}(\mathbf{s})$ as:
$p(\mathbf{x})=p_{\theta}(\mathbf{s})|J_{\mathbf{f}^{-1}}(\mathbf{x})|.$ (4)
[…]
$p(\mathbf{x})=p_{\hat{\theta}}(\mathbf{s})|J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x})|.$ (5)
We define the indeterminacy function
$\mathbf{h}:=\widehat{\mathbf{f}}^{-1}\circ\mathbf{f},$ (6)
which "quantifies" how different the two ICA solutions are. By the multivariate chain rule, the following relation among Jacobian matrices holds:
$J_{\mathbf{f}}=J_{\widehat{\mathbf{f}}}J_{\mathbf{h}}.$ (7)
[…]
4.1 Identifiability from second order derivatives of the log-likelihood
In this section, we present our main theoretical result and the intuitions behind it. Our argument for identifiability relies on the analysis of the Hessian of the log-likelihood of $\mathbf{X}^{i}$ for all environments. We consider the case where $\mathbf{f}^{-1}(\mathbf{x})=\mathbf{s}=\mu_{\mathbf{S}}$ (by construction, there is a unique corresponding $\hat{\mathbf{s}}=\hat{\mathbf{f}}^{-1}(\mathbf{x})$). We partition the set of $e$ auxiliary environments into two groups $I_{1}=\{1,...,e_{1}\}$ and $I_{2}=\{e_{1}+1,...,e_{1}+e_{2}\}$, where $e=e_{1}+e_{2}$. Then, we define the following quantities:
$\begin{split}&\Omega_{1}:=\sum_{i\in I_{1}}D^{2}_{\mathbf{s}}\log p_{\theta}(\mathbf{s})-D^{2}_{\mathbf{s}}\log p^{i}_{\theta}(\mathbf{s})\\ &\Omega_{2}:=\sum_{i\in I_{2}}D^{2}_{\mathbf{s}}\log p_{\theta}(\mathbf{s})-D^{2}_{\mathbf{s}}\log p^{i}_{\theta}(\mathbf{s}),\end{split}$ (8)
where $D^{2}$ denotes the differential operator that returns the Hessian matrix. Similarly, we define $\widehat{\Omega}_{1},\widehat{\Omega}_{2}$ by replacing $\theta$ with $\hat{\theta}$. The introduction of $\Omega_{l},\widehat{\Omega}_{l}$, $l=1,2$, is instrumental for the next result.
Lemma 1.
[…]
$\sum_{i\in I_{1}}D_{\mathbf{x}}^{2}\log p(\mathbf{x})-D_{\mathbf{x}}^{2}\log p^{i}(\mathbf{x})=J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}\Omega_{1}J_{\mathbf{f}^{-1}}(\mathbf{x})=J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x})^{T}\widehat{\Omega}_{1}J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x})$ (9)
$\sum_{i\in I_{2}}D_{\mathbf{x}}^{2}\log p(\mathbf{x})-D_{\mathbf{x}}^{2}\log p^{i}(\mathbf{x})=J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}\Omega_{2}J_{\mathbf{f}^{-1}}(\mathbf{x})=J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x})^{T}\widehat{\Omega}_{2}J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x})$ (10)
The proof is derived by direct computation and can be found in Section˜A.2. We can intuitively illustrate how the identifiability of the Jacobian’s support follows. A first remark is that the $\Omega_{l},\widehat{\Omega}_{l}$ matrices are diagonal. That is because, for a vector of mutually independent random variables, the Hessian of the log-density is diagonal (see Section˜D.3 for details about it). Second, by the chain rule, Equations˜9 and 10 imply $J_{\mathbf{h}}(\mathbf{s})^{T}\widehat{\Omega}_{l}J_{\mathbf{h}}(\mathbf{s})=\Omega_{l}$ for $l=1,2$, from which
$J_{\mathbf{h}}(\mathbf{s})^{-1}\widehat{\Omega}_{1}^{-1}\widehat{\Omega}_{2}J_{\mathbf{h}}(\mathbf{s})=\Omega_{1}^{-1}\Omega_{2}.$ (11)
This means that $J_{\mathbf{h}}(\mathbf{s})$ maps one diagonal matrix to another: if the eigenvalues of $\widehat{\Omega}_{1}^{-1}\widehat{\Omega}_{2}$ are distinct, that is enough to force $J_{\mathbf{h}}(\mathbf{s})$ to a scaled permutation, which is exactly our goal. This sketched argument is key to understanding how Equations˜9 and 10 provide enough constraints to identify the support of $J_{\mathbf{f}^{-1}}$. Clearly, this discussion implicitly requires that $\Omega_{l}$ and $\widehat{\Omega}_{l}$ are full rank. This can be achieved under the following conditions over the rescaling matrices $L_{i}=\mathrm{diag}(\lambda_{1}^{i},\dots,\lambda_{d}^{i})$ that define the multiple environments.
Assumption 5 (Sufficient variability).
For each $j\in[d]$:
[…]
The assumption basically requires that there is sufficient variability between the different environments. Similar requirements of sufficient variability are ubiquitous in the nonlinear ICA literature (e.g. Hyvärinen & Morioka (2016); Khemakhem et al. (2020b); Lachapelle et al. (2022)). Intuitively speaking, ˜5 is satisfied when, for each of the two groups of environments ($1,...,e_{1}$ and $e_{1}+1,...,e_{1}+e_{2}$), each source $S_{j}$ is subject to rescaling. To see that, consider the LHS of the first equation: $\lambda_{j}^{i}=1$ for each $i=1,...,e_{1}$ corresponds to the case when the variable $S_{j}$ is never subject to rescaling in any of the environments, and indeed yields a violation of the assumption. Note that even if $S_{j}$ is subject to rescaling for some index $i$, the values of $(\lambda_{j}^{i})_{i\in[e_{1}]}$ can always be tuned such that the assumption is violated; however, this corresponds to pathological choices of the rescaling coefficients, which never occur in general (shown in Proposition˜3 in the appendix).
Next, we are ready to state our main identifiability result.
Theorem 1.
[…]
Theorem˜1 assumes that the elements in the set $\{(\Omega_{1}^{-1}\Omega_{2})_{ii}\}_{i=1}^{d}$ are pairwise distinct. This requirement excludes pathological choices of the coefficients of the rescaling matrices $L_{i}$ that define the multiple environments, and it is generically satisfied (Proposition˜4 in the appendix).
Proof sketch (full proof in Section˜A.4).
By Lemma˜1 we have
$M^{T}\Omega_{l}M=\widehat{\Omega}_{l},\quad l=1,2,$ (12)
[…]
Identifiability from two auxiliary environments.
The theorem tells that, given that we have access to two groups of auxiliary environments, both inducing changes in the variance of all sources, at the mean of the sources the ground truth and the alternative models are equivalent up to rescaling. This constrains the support of $J_{\widehat{\mathbf{f}}^{-1}}$ of the alternative model to be equal to that of $J_{\mathbf{f}^{-1}}$, which is enough to guarantee identifiability of the causal graph. It is interesting to discuss the theorem when $e_{1}+e_{2}=2$, showing that the above result demonstrates identifiability with as few as two additional environments. In this setting, if $L_{1}=\operatorname{diag}(\lambda_{j}^{1})_{j=1}^{d}$ and $L_{2}=\operatorname{diag}(\lambda_{j}^{2})_{j=1}^{d}$ with $\lambda_{j}^{1},\lambda_{j}^{2}\neq 1$ for each $j\in[d]$, then we have two extra environments where the variance of all the sources is affected by rescaling. This is sufficient to guarantee that the assumptions of Theorem˜1 are met. An important consequence is that the number of required environments does not scale with the number of nodes in the graph, in contrast with similar findings for nonlinear ICA identifiability. As long as there is sufficient variability in the sources of two environments (relative to the base model), we are always guaranteed that the causal graph can be recovered.
Theorem˜1 beyond Gaussianity.
Theorem˜1 inherits the assumption of Gaussianity from Lemma˜1; here, we briefly discuss potential ways to relax it. At a general point $\mathbf{x}=\mathbf{f}(\mathbf{s})$ the Hessian of the log-likelihood is equal to
$J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}D^{2}_{\mathbf{\mathbf{s}}}\log p^{i}(\mathbf{s})J_{\mathbf{f}^{-1}}(\mathbf{x})+D_{\mathbf{x}}^{2}\log|J_{\mathbf{f}^{-1}}(\mathbf{x})|+\sum_{j=1}^{d}\partial s_{k}\log p^{i}(s_{j})D^{2}\mathbf{f}^{-1}_{j}(\mathbf{x}).$
The log-determinant term cancels by taking the difference between environments. To recover Equations˜9 and 10 in Lemma˜1, we note that the summation of second-order derivatives vanishes when $\nabla\log p^{i}(\mathbf{s})=0$, namely at the mean of the Gaussian sources. However, this can hold for any source distribution that has at least one point where the gradient is zero, a remark that naturally extends Lemma˜1 (and hence, Theorem˜1) to a larger class of causal models. Moreover, from a practical perspective, even if the gradient of the log-likelihood of the sources does not vanish, Lemma˜1 is approximately true when the gradient is sufficiently small. This can occur, e.g., for heavy-tailed distributions. This analysis should convince that Gaussianity is a sufficient but not necessary requirement, and hopefully inspire future research to extend our identifiability results.
Next, we support the conclusions of our theory with experiments.
5 Empirical results
In this section, we report and analyse empirical results that validate our theory. Our experiments on synthetic data show that if the assumptions of Theorem˜1 hold, the causal direction can be recovered from the data. We focus on bivariate graphs, commonly adopted as the easiest yet non-trivial setting for testing identifiability (e.g., Hoyer et al. (2008); Zhang & Hyvärinen (2009); Immer et al. (2022)).
5.1 Synthetic data generation
We generate synthetic data from bivariate causal models with independent noise terms, sampled from a normal distribution with unit mean and covariance entries uniformly drawn between $[1,1.5]$. Given the variables $x_{1},x_{2}$ and the graph $x_{1}\to x_{2}$ we consider the following causal mechanisms that comply with the assumptions of Theorem˜1: (i) $x_{2}:=s_{1}^{2}\operatorname{arctan}(s_{2})+s_{2}^{3}$ (ii) $x_{2}:=s_{1}^{2}s_{2}+\operatorname{arctan}(s_{2})$ (iii) $x_{2}:=s_{1}^{2}+\operatorname{arctan}(s_{1})s_{2}+s_{1}s_{2}^{3}$. Note that any of these models can not be reparametrized to a post nonlinear or location scale noise model, which are the most general SCMs identifiable from pure observations (Zhang & Hyvärinen, 2009; Immer et al., 2022). Additionally, we consider data from a linear Gaussian model, notably non-identifiable. We run experiments on datasets with $\{3,6,9\}$ environments. For each environment, we generate $2000$ observations. In Section˜C.4, we discuss experiments with non-Gaussian independent sources. Interestingly, these additional results seem to support our hypothesis that Theorem˜1 could be extended to other source distributions.
5.2 Analysis of the experimental results
In this section, we analyse the empirical results. First, we introduce an algorithm for inferring the Jacobian support that leverages our theory.
Data: $\widehat{X}\in\mathbb{R}^{k\times n\times d}$
// $\forall$ env: $n$ d-dimensional observations.
    $I_{1},I_{2}\subset[k]$
// Set of indices splitting the environments in two groups
Result: Estimate of $\operatorname{supp}J_{\mathbf{f}^{-1}}$
$\widehat{S}\leftarrow\operatorname{score\_estimate}(\widehat{X})\in\mathbb{R}^{k\times n\times d}$
$\widehat{H}\leftarrow\operatorname{hess\_estimate}(\widehat{X})\in\mathbb{R}^{k\times n\times d\times d}$
// For each environment $e$, find the sample corresponding to the mean of the source
for e=1,,ke=1,...,k do
 $m_{e}\leftarrow i$ s.t. $\mathbf{f}^{-1}(\widehat{X}[e,i])\approx\mu_{\mathbf{S}}$
end for
// Difference of Hessians at the mean (i.e. Equations˜9 and 10)
$\widehat{H}_{\textnormal{diffs}}\leftarrow 0\in\mathbb{R}^{2\times d\times d}$
for $\ell=1,2$ do
 for $e\in I_{\ell}$ do
    $\Delta_{H}=\widehat{H}[0,m_{1}]-\widehat{H}[e,m_{e}]$
    // $m_{1}$ is the index for the base environment
    $\widehat{H}_{\textnormal{diffs}}[\ell]\leftarrow\widehat{H}_{\textnormal{diffs}}[\ell]+\Delta_{H}$.
   end for
end for
$M\leftarrow\widehat{H}_{\textnormal{diffs}}^{-1}[1]\widehat{H}_{\textnormal{diffs}}[2]\approx J_{\mathbf{f}}\Omega_{1}^{-1}\Omega_{2}J_{\mathbf{f}^{-1}}$
// $H_{\textnormal{diffs}}[\ell]\approx J_{\mathbf{f}^{-1}}^{T}\Omega_{\ell}J_{\mathbf{f}^{-1}}$, by Equations˜9 and 10
$\widehat{J}_{\mathbf{f}^{-1}}\leftarrow\textnormal{diagonalize}(M)\approx J_{\mathbf{f}^{-1}}DP$
return $\operatorname{supp}\left(\widehat{J}_{\mathbf{f}^{-1}}P^{-1}\right)$
// $P$ can be found using the acyclicity of the causal graph.
Algorithm 1 Estimating $\operatorname{supp}J_{\mathbf{f}^{-1}}$ from the data (algorithm sketch)
Algorithm.
The simplified pseudocode is found in Algorithm˜1 (a detailed version is presented in Section˜C.3). The steps in our procedure closely follow the proof of Theorem˜1: this approach to algorithmic design is not necessarily the best, which is why we highlight that our method is not within our main contributions. For a single inference, the input is the data tensor $\widehat{X}\in\mathbb{R}^{k\times n\times d}$: for each environment from $1$ to $k$ it consists of a dataset with $n$ observations of $d$ causal variables. Additionally, we are given the sets $I_{1},I_{2}\subset[k]$ of indices that split the auxiliary environments into two groups, as required by our theory. The first environment is taken as the base one. We have two steps where statistical estimation is involved: (i) For each environment, the gradient and the Hessian of the log-likelihood are approximated via the Stein gradient estimator, introduced in Li & Turner (2018) and popularized in causal discovery by Rolland et al. (2022); Montagna et al. (2023b); (ii) For each environment $i\in[k]$, we need to find the observation $j\in[n]$ such that $\mathbf{f}^{-1}(\widehat{X}[i,j])\approx\mu_{\mathbf{S}}$, that is, the data point generated mixing the source vector at the mean. Fortunately, this can be consistently estimated from the score $\nabla\log p_{\mathbf{x}}$, as we demonstrate in Proposition˜2 in the appendix. These two steps are achieved by Algorithm˜1 at the end of the first for loop. At this stage, all statistical quantities have been estimated: we note that, being the Stein estimator consistent, the algorithm is correct in the infinite sample limit. In the second for loop, we take the points at the estimated mean that we previously found, and compute the difference of the Hessians between the base and auxiliary environments: this exactly mirrors the first equality in Equations˜9 and 10 of Lemma˜1. Next, in the algorithm’s notation, we compute
$M:=\widehat{H}^{-1}_{\textnormal{diffs}}[1]\widehat{H}_{\textnormal{diffs}}[2]\approx J_{\mathbf{f}}\Omega_{1}^{-1}\Omega_{2}J_{\mathbf{f}^{-1}}.$ (13)
Then, we solve the linear system $\widehat{H}_{\textnormal{diffs}}[1]M=\widehat{H}_{\textnormal{diffs}}[2]$ to find $M$. In the infinite samples limit Equation˜13 is a precise equality, such that $M$ and $\Omega_{1}^{-1}\Omega_{2}$ are similar: diagonalizing $M$ we find $J_{\mathbf{f}^{-1}}$ up to a scaled permutation. The permutation indeterminacy is removed leveraging the assumption that the causal graph is acyclic via standard arguments (see Shimizu et al. (2006) and Reizinger et al. (2023)). Finally, the algorithm returns the estimated support of the inverse Jacobian.
Analysis of the experiments.
In Figure˜1 we illustrate the empirical performance of our method on several synthetic datasets generated from a bivariate causal model. We consider SCMs with the arbitrary nonlinear mechanisms (i), (ii), (iii) described in Section˜5.1, and linear Gaussian models; as a sanity check, we also experiment on nonlinear additive noise models (ANM), post-nonlinear models (PNL), and location scale noise models (LSNM), which are all the nonlinear SCMs where identifiability can be achieved from observational data (see Section˜C.2 for details). All datasets are generated under the assumption that a causal effect exists (i.e., the ground truth graphs always have one arrow). We measure the errors through the structural hamming distance (SHD). This is equivalent to the number of edge additions, removals, or direction flips that are required to recover the ground truth graph from the estimated one: SHD=0 corresponds to correct inference, SHD=1 to an error. For each experimental configuration, consisting of function type and number of environments, we consider $50$ seeds over which we compute the empirical mean and deviation of the SHD. The results are in line with our theory: we see that for the three models with arbitrary mechanisms, and the linear Gaussian SCM (all non-identifiable from pure observations), the average SHD is close to 0, which is especially evident when we do inference with only $3$ environments. Interestingly, we see that adding environments doesn’t always have a beneficial effect. This is not surprising, as we showed that two auxiliary environments are sufficient for inference. The method can also infer the causal direction for the ANM, PNL, and LSNM. We conclude that the empirical outcomes support our theory.
Figure 1: Average SHD (0 is best, 11 is worst) achieved by Algorithm˜1 over 5050 seeds on binary graphs. When the assumptions of Theorem˜1 are satisfied, the method can appropriately infer the causal direction, both in the observationally identifiable setting (nonlinear ANM, PNL, LSNM) and the observationally non-identifiable one (linear Gaussian model and the three SCMs with arbitrary nonlinearity). The number of environments does not have a notable effect on the accuracy.
6 Conclusion
We demonstrated that the causal graph of a structural causal model with arbitrary nonlinear mechanisms is identifiable; surprisingly, this can be achieved given the auxiliary information of only two (sufficiently different) environments. Our main assumption is the Gaussianity of the noise terms, for which, however, we discuss potential relaxations. Our findings extend on the well-known duality between ICA and causal discovery: the first problem concerns the identifiability of the independent sources at each point, whereas causality only needs to access the support of the Jacobian mixing function at a single point, when faithfulness is satisfied. The exciting consequence of this asymmetry is that while ICA identifiability requires a number of environments that grows linearly with the number of sources, for causal discovery, a constant number is sufficient: this makes our theoretical results appealing even in high dimensions. We hope that our work inspires novel identifiability theory beyond the Gaussianity constraint. Moreover, in light of our results, finding an efficient and effective algorithm for causal discovery with multiple environments and in high dimensions is a promising research direction.
Appendix A Proof of the theoretical results
A.1 Preliminary theoretical results
In this section, we collect the theoretical results useful for the proof of Theorem˜1.
Lemma 2 (Full rank of $\Omega_{l}$ under rescalings).
Assume Gaussian sources $\mathbf{S}$ with independent coordinates, and environments generated by rescalings $\mathbf{S}^{i}=L_{i}\mathbf{S}$ with $L_{i}=\mathrm{diag}(\lambda_{1}^{i},\dots,\lambda_{d}^{i})$ and $\lambda_{j}^{i}\neq 0$. For $l\in\{1,2\}$ define the index sets $I_{1}=\{1,\dots,e_{1}\}$ and $I_{2}=\{e_{1}+1,\dots,e_{1}+e_{2}\}$, and recall
$\Omega_{l}\ :=\ \sum_{i\in I_{l}}\Big(D_{s}^{2}\log p_{\theta}(\mathbf{s})\ -\ D_{s}^{2}\log p_{\theta}^{i}(\mathbf{s})\Big),$
evaluated at the same $\mathbf{s}$. Then each $\Omega_{l}$ is diagonal with entries
$(\Omega_{l})_{jj}\ =\ \frac{1}{\sigma_{j}^{2}}\left(\sum_{i\in I_{l}}\frac{1}{(\lambda_{j}^{i})^{2}}\ -\ |I_{l}|\right),$
and therefore
$\Omega_{l}\text{ is full rank }\iff\forall j\in[d]:\ \sum_{i\in I_{l}}\frac{1}{(\lambda_{j}^{i})^{2}}\ \neq\ |I_{l}|.$
Proof.
For a univariate Gaussian, $D_{s_{j}}^{2}\log p(s_{j})=-1/\sigma_{j}^{2}$. In environment $i$ we have $S_{j}^{i}=\lambda_{j}^{i}S_{j}$, so $S_{j}^{i}$ has variance $(\lambda_{j}^{i}\sigma_{j})^{2}$, hence $D_{s_{j}}^{2}\log p^{i}(s_{j})=-1/(\lambda_{j}^{i}\sigma_{j})^{2}$. Thus
$\big(D_{s_{j}}^{2}\log p(s_{j})-D_{s_{j}}^{2}\log p^{i}(s_{j})\big)=\frac{1}{\sigma_{j}^{2}}\Big(\frac{1}{(\lambda_{j}^{i})^{2}}-1\Big).$
Summing over $i\in I_{l}$ gives the stated diagonal form. A diagonal matrix is full rank iff none of its diagonal entries is zero, which yields the equivalence. ∎
Lemma 3.
$\Omega_{l}$ is invertible implies $\widehat{\Omega}_{l}$ invertible.
Proof.
By Lemma˜1, for $l=1,2$, we have:
$J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}\Omega_{l}J_{\mathbf{f}^{-1}}(\mathbf{x})=J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x})^{T}\widehat{\Omega}_{l}J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x}).$
Under ˜5, by Lemma˜2 the LHS is a product of full rank matrices, and so is full rank; so must be the RHS. Given that $\operatorname{rank}(AB)\leq\min\left(\operatorname{rank}(A),\operatorname{rank}(B)\right)$ (for generic matrices $A,B$) we conclude that $\widehat{\Omega}_{l}$ is also full rank. ∎
A.2 Proof of Lemma 1
We report the content of Lemma˜1, followed by its proof. See 1
Proof.
By direct computation, it can be verified that for each $i=0,...,e_{1}+e_{2}$, we have:
$\begin{split}D^{2}_{\mathbf{x}}\log p^{i}(\mathbf{x})&=D^{2}_{\mathbf{x}}\log|J_{\mathbf{f}^{-1}}(\mathbf{x})|+J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}D^{2}_{\mathbf{s}}\log p^{i}_{\theta}(\mathbf{s})J_{\mathbf{f}^{-1}}(\mathbf{x})\\ &+\sum_{k=1}^{d}\partial_{s_{k}}\log p^{i}_{\theta}(s_{k})D^{2}_{\mathbf{x}}\mathbf{f}^{-1}_{k}(\mathbf{x}).\end{split}$ (14)
Given $\mathbf{s}=\mu_{\mathbf{S}}$, ˜4 of Gaussianity, together with the fact that $\mathbf{S}^{i}=L_{i}\mathbf{S}$ for some diagonal $L_{i}$, imply $\partial_{s_{k}}\log p^{i}_{\theta}(s_{k})=0$ for all $k$. Then, the summation vanishes. It follows that, for all environments $i=1,...,e_{1}+e_{2}$:
$D^{2}_{\mathbf{x}}\log p(\mathbf{x})-D^{2}_{\mathbf{x}}\log p^{i}(\mathbf{x})=J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}\left(D^{2}_{\mathbf{s}}\log p^{i}_{\theta}(\mathbf{s})-D^{2}_{\mathbf{s}}\log p^{i}_{\theta}(\mathbf{s})\right)J_{\mathbf{f}^{-1}}(\mathbf{x}).$
The same results hold if we replace $\mathbf{f}$ with $\widehat{\mathbf{f}}$ and $\theta$ with $\hat{\theta}$. Then, Equation˜9 follows summing the above over all $i=1,...,e_{1}$, and Equation˜10 follows summing over $i=e_{1}+1,...,e_{1}+e_{2}$. ∎
A.3 Identifiability of the mean of the sources
In this section, we show that under the assumptions of Theorem˜1, the mean $\mu_{\mathbf{S}}$ of the sources is identifiable.
Proposition 2 (Identifiability of the sources mean).
For each $i=1..,e_{1}+e_{2}$, suppose the diagonal entries of the rescaling matrices $L_{i}$ generating the environments are randomly drawn from a joint distribution that is absolutely continuous with respect to the Lebesgue measure on $(\mathbb{R}\setminus{0})^{d(e_{1}+e_{2})}$. Then, the following is verified with probability one over the samples $\{L_{i}\}_{i=1}^{e_{1}+e_{2}}$:
$\sum_{i=1}^{k}\nabla\log p(\mathbf{x})-\nabla\log p^{i}(\mathbf{x})=0\iff\mathbf{s}=\mathbf{f}^{-1}(\mathbf{x})=\mu_{\mathbf{S}}.$ (15)
We introduce two lemmas instrumental to the proof of the proposition.
Lemma 4.
Consider the base ICA model of Equation˜2, and let i=1,,ki=1,...,k be the index denoting an auxiliary environment (Definition˜3). Let Assumptions 1-4 to be satisfied. Given $\mathbf{x}=\mathbf{f}(\mathbf{s})$ such that $J_{\mathbf{f}^{-1}}(\mathbf{x})$ is full rank, for each $k\leq e_{1}+e_{2}$:
$\sum_{i=1}^{k}\nabla\log p(\mathbf{x})-\nabla\log p^{i}(\mathbf{x})=0\iff\sum_{i=1}^{k}\nabla\log p(\mathbf{s})-\nabla\log p^{i}(\mathbf{s})=0$ (16)
Proof.
By the change of variable formula for densities, we obtain the score of $\mathbf{x}$ for a generic environment i=0,,ki=0,...,k (as usual, $p=p^{0}$):
$\nabla\log p^{i}(\mathbf{x})=J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}\nabla\log p^{i}(\mathbf{s})+\nabla\log|J_{\mathbf{f}^{-1}}(\mathbf{x})|.$
Then, for each i=1,,ki=1,...,k:
$\nabla\log p(\mathbf{x})-\nabla\log p^{i}(\mathbf{x})=J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}\left[\nabla\log p(\mathbf{s})-\nabla\log p^{i}(\mathbf{s})\right].$
Taking the summation:
$\sum_{i=1}^{k}\nabla\log p(\mathbf{x})-\nabla\log p^{i}(\mathbf{x})=\sum_{i=1}^{k}J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}\left[\nabla\log p(\mathbf{s})-\nabla\log p^{i}(\mathbf{s})\right].$
From the above equation, the right-to-left implication trivially holds. Considering the other direction we have:
$\sum_{i=1}^{k}\nabla\log p(\mathbf{x})-\nabla\log p^{i}(\mathbf{x})=0\implies\sum_{i=1}^{k}J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}\left[\nabla\log p(\mathbf{s})-\nabla\log p^{i}(\mathbf{s})\right]=0.$
Being the Jacobian of the inverse mixing function a full rank matrix, its null space is the zero vector, which implies:
$\sum_{i=1}^{k}\nabla\log p(\mathbf{s})-\nabla\log p^{i}(\mathbf{s})=0.$
Lemma 5.
Consider the base ICA model $\mathbf{X}=\mathbf{f}(\mathbf{S})$ of Equation˜2. Let i=1,,ki=1,...,k be the index of the auxiliary environment $\mathbf{X}^{i}=\mathbf{f}(\mathbf{S}^{i})$, with $\mathbf{S}^{i}=L_{i}\mathbf{S}$, $L_{i}=\mathrm{diag}(\lambda_{1}^{i},\dots,\lambda_{d}^{i})$, and $\lambda_{j}^{i}\neq 0$. Let Assumptions 1 and 4 be satisfied. Assume the joint law of $\{\lambda_{j}^{i}:j=1,\dots,d,\ i=1,\dots,k\}$ is absolutely continuous with respect to Lebesgue measure on $(\mathbb{R}\setminus\{0\})^{dk}$. Then, for each $k\leq e_{1}+e_{2}$, the following holds with probability one over $\{L_{i}\}_{i=1}^{k}$ samples:
$\sum_{i=1}^{k}\nabla\log p(\mathbf{s})-\nabla\log p^{i}(\mathbf{s})=0\iff\mathbf{s}=\mathbf{f}^{-1}(\mathbf{x})=\mu_{\mathbf{S}}.$ (17)
Proof.
The backward direction is immediate, due to the Gaussianity assumption. Let’s focus on the forward implication.
$\sum_{i=1}^{k}\nabla\log p(\mathbf{s})-\nabla\log p^{i}(\mathbf{s})=0\iff\sum_{i=1}^{k}\partial_{s_{j}}\log p(s_{j})-\partial_{s_{j}}\log p^{i}(s_{j})=0,\hskip 5.0pt\forall j=1,...,d.$
We denote with $\mu_{j},\sigma_{j}^{2}$ respectively the mean and variance of $S_{j}$, and define $\lambda^{0}_{j}:=1$. For each i=0,,ki=0,...,k we have:
$\partial_{s_{j}}\log p^{i}(s_{j})=\frac{\mu_{j}-s_{j}}{(\lambda_{j}^{i}\sigma_{j})^{2}}.$
Then:
$\sum_{i=1}^{k}\partial_{s_{j}}\log p(s_{j})-\partial_{s_{j}}\log p^{i}(s_{j})=\frac{\mu_{j}-s_{j}}{\sigma_{j}^{2}}\left(k-\sum_{i=1}^{k}\frac{1}{(\lambda_{j}^{i})^{2}}\right).$
Therefore, the sum vanishes if and only if for every $j$, either $s_{j}=\mu_{j}$ or $\sum_{i=1}^{k}(\lambda_{j}^{i})^{-2}=k.$ By Proposition˜3, $\sum_{i=1}^{k}(\lambda_{j}^{i})^{-2}=k$ occurs with probability zero, and thus the claim is verified.
We are ready to prove the proposition.
Proof of Proposition˜2.
By Lemma˜4 we have that for each $k\leq e_{1}+e_{2}$:
$\sum_{i=1}^{k}\nabla\log p(\mathbf{x})-\nabla\log p^{i}(\mathbf{x})=0\iff\sum_{i=1}^{k}\nabla\log p(\mathbf{s})-\nabla\log p^{i}(\mathbf{s})=0$
Then, the result follows by application of Lemma˜5. ∎
A.4 Proof of Theorem 1
We repropose the statement of Theorem˜1, followed by a detailed proof. See 1
Proof.
By Lemma˜1, for $l=1,2$ we have:
$J_{\mathbf{f}^{-1}}(\mathbf{x})^{T}\Omega_{l}J_{\mathbf{f}^{-1}}(\mathbf{x})=J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x})^{T}\widehat{\Omega}_{l}J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x}),$
which implies
$M^{T}\Omega_{l}M=\widehat{\Omega}_{l},\quad l=1,2,$ (18)
[…]
$A=M^{-1}BM,$ (19)
which implies that $A$ and $B$ are similar, implying that they have the same set of eigenvalues. Take $\lambda,\mathbf{v}$ eigenvectors of $A$. Then, the following chain of implication holds:
$A\mathbf{v}=\lambda\mathbf{v}\iff MA\mathbf{v}=\lambda M\mathbf{v}\iff BM\mathbf{v}=\lambda M\mathbf{v},$ (20)
where the last step follows from Equation˜19. So, $M$ is mapping from eigenvectors of $A$ to eigenvectors of $B$. The next step is showing that each eigenspace of $A$ and $B$ is always spanned by one vector in the standard basis. As a preliminary step, we show that the diagonal elements of $A$ are pairwise distinct: first, by similarity, we have that $A$ and $B$ have the same eigenvalues. Being both matrices diagonal, the eigenvalues are the diagonal elements. Then:
$A_{ii}=\frac{(\widehat{\Omega}_{1})_{ii}}{(\widehat{\Omega}_{2})_{ii}}=\frac{(\Omega_{1})_{jj}}{(\Omega_{2})_{jj}}=B_{jj},\hskip 10.00002pti,j=1,...,d.$ (21)
By assumption, we have that the elements in the set $\{\frac{(\Omega_{1})_{\ell\ell}}{(\Omega_{2})_{\ell\ell}}\}_{\ell\in[d]}$ are pairwise distinct. The above equation implies the same for the set $\{\frac{(\widehat{\Omega}_{1})_{\ell\ell}}{(\widehat{\Omega}_{2})_{\ell\ell}}\}_{\ell\in[d]}$, i.e., for each i=1,,di=1,...,d:
$A_{ii}\neq A_{jj},\hskip 10.00002pt\forall j=1,...,d,j\neq i.$ (22)
Now consider the eigenvalue $\lambda$ of $A$: we show that the associated eigenspace is equal to the span of a single vector in the standard basis. Being $A$ diagonal, there is i=1,,di=1,...,d such that $\lambda=A_{ii}$. Consider the eigenvector $\mathbf{v}=(v_{1},...,v_{d})$ such that:
$A\mathbf{v}=\lambda\mathbf{v}=A_{ii}\mathbf{v}.$ (23)
Being $A$ diagonal, for each j=1,,dj=1,...,d, component-wise we have:
$(A\mathbf{v})_{j}=A_{jj}v_{j}.$ (24)
Equations˜23 and 24 together imply:
$A_{ii}v_{j}=A_{jj}v_{j}\iff(A_{ii}-A_{jj})v_{j}=0,\hskip 10.00002pt\forall j=1,...,d.$
By Equation˜22, for $i\neq j$, $A_{ii}\neq A_{jj}$, meaning that $v_{j}=0$. Then, $\mathbf{v}$ eigenvector of $A$ must be aligned with the basis vector $\mathbf{e}_{i}$:
$E_{\lambda}(A)=\operatorname{span}\{\mathbf{e}_{i}\}.$ (25)
With analogous computations, we find:
$E_{\lambda}(B)=\operatorname{span}\{\mathbf{e}_{j}\},$ (26)
with $\mathbf{e}_{j}$ potentially different from $\mathbf{e}_{i}$. Given that by Equation˜20 we have $ME_{\lambda}(A)=E_{\lambda}(B)$, the last two equations imply
$M\operatorname{span}\{\mathbf{e}_{i}\}=\operatorname{span}\{\mathbf{e}_{j}\},\hskip 10.00002ptM=J_{\mathbf{h}}(\mathbf{s}).$
We conclude that $J_{\mathbf{h}}(\mathbf{s})$ maps one vector in the standard basis to another (up to rescaling), proving that $J_{\mathbf{h}}(\mathbf{s})=DP$ with $D$ invertible diagonal and $P$ permutation. We recall that by Equation˜7 we have $J_{\mathbf{f}}=J_{\widehat{\mathbf{f}}}J_{\mathbf{h}}$, s.t.
$J_{\mathbf{f}^{-1}}(\mathbf{x})=P^{T}D^{-1}J_{\widehat{\mathbf{f}}^{-1}}(\mathbf{x}).$
By Lemma 1 in Reizinger et al. (2023), the permutation indeterminacy can be uniquely determined and thus removed. Given that by ˜3 the Jacobian of $J_{\mathbf{f}^{-1}}(\mathbf{x})$ is faithful to the causal graph, the claim is verified.
Appendix B Independent component analysis
In this section, we present a primer on the problem of Independent Component Analysis (ICA), based on the content of Section 2 in Buchholz et al. (2022). ICA seeks to recover latent sources from their observed mixtures. We assume a hidden random vector $\mathbf{S}\in\mathbb{R}^{d}$ with independent coordinates and observations generated by
$\mathbf{X}\;=\;\mathbf{f}(\mathbf{S}),\qquad p(\mathbf{s})\;=\;\prod_{i=1}^{d}p_{i}(s_{i}),$ (27)
where $\mathbf{f}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a diffeomorphism. The goal of ICA is to find an unmixing map $\widehat{\mathbf{f}}^{-1}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that the components of $\widehat{\mathbf{f}}^{-1}(\mathbf{X})$ are independent—ideally achieving blind source separation (BSS), meaning $\widehat{\mathbf{f}}^{-1}\approx\mathbf{f}^{-1}$ up to standard symmetries. Informally, for $\hat{\mathbf{s}}=\widehat{\mathbf{f}}^{-1}(\mathbf{x})$, we call $\widehat{\mathbf{f}}$ an ICA solution when
$\widehat{\mathbf{f}}(\hat{\mathbf{s}})\stackrel{{\scriptstyle D}}{{=}}\mathbf{f}(\mathbf{s})$
(equality is in distribution). In general, we would like an ICA solution to be as close as possible to the real function $\mathbf{f}$. To formalize this concept, known as identifiability, let $\mathcal{F}(\mathcal{A},\mathcal{B})$ be a class of invertible maps $\mathcal{A}\to\mathcal{B}$ (assumed diffeomorphisms) and let $\mathcal{P}\subset\mathcal{M}_{1}(\mathbb{R})^{\otimes d}$ be a family of product measures. Let $\mathcal{S}$ denote the group of admissible symmetries (e.g., permutations and coordinate-wise rescalings) up to which we agree to identify sources.
Definition 5 (Identifiability).
ICA in $(\mathcal{F},\mathcal{P})$ is identifiable up to $\mathcal{S}$ if, for any $\mathbf{f},\widehat{\mathbf{f}}\in\mathcal{F}$ and $P,\hat{P}\in\mathcal{P}$,
[…]
implies the existence of $\mathbf{h}\in\mathcal{S}$ such that $\mathbf{h}=\widehat{\mathbf{f}}^{-1}\!\circ\mathbf{f}$ on the support of $P$.
In general (i.e., for $(\mathcal{F},\mathcal{P})$ arbitrarily large), the ICA problem is not identifiable for reasonable $\mathcal{S}$. Notable example comes from the Darmois construction or constructions based on measure-preserving transformations. Several results in the literature have studied which conditions on $(\mathcal{F},\mathcal{P})$ can help identifiability. Most notably, Buchholz et al. (2022) shows that when $\mathcal{F}$ represents the class of conformal maps, identifiability is guaranteed up to trivial indeterminacies. If heterogeneous data are considered (e.g., in the multi-environment setting of this paper), identifiability was shown in the general case (Hyvärinen & Morioka, 2016).
Appendix C Experiments appendix
C.1 Computational resources
All experiments have been run on a personal laptop, a Lenovo ThinkPad T14 Gen 5, for a run time of approximately $6$ hours.
C.2 Structural causal model identifiability from observational data
Without sufficiently restrictive modeling assumptions, causal discovery is ill-posed: the distribution of the data is compatible with many distinct graphs that define an equivalence class, the most one can hope to identify in the general case with i.i.d. observations. Unique graph recovery requires restrictions on the class of functional mechanisms and noise distributions of the underlying causal model: in what follows, we briefly introduce the four classes of causal models that are known to be identifiable. We always assume that the underlying graph is a DAG.
Linear Non-Gaussian Model (LiNGAM).
A linear SCM over $\mathbf{X}\in\mathbb{R}^{d}$ is defined by
$\mathbf{X}=B\mathbf{X}+\mathbf{S},$ (29)
where $B\in\mathbb{R}^{d\times d}$ collects the coefficients expressing each $X_{i}$ as a linear function of its parents plus a disturbance $S_{i}$. With mutually independent, non-Gaussian noise terms, the model is identifiable; this is known as the Linear Non-Gaussian Acyclic Model (LiNGAM) (Shimizu et al., 2006).
Additive Noise Model (ANM).
An Additive Noise Model (ANM) (Hoyer et al., 2008; Peters et al., 2014) defines each causal variable as a function of (potentially) nonlinear mechanisms and an additive noise contribution:
$X_{i}\coloneqq f_{i}(\operatorname{PA}_{i})+S_{i},\quad i=1,\ldots,d.$ (30)
The noise terms are required to be mutually independent.
Post-Nonlinear Model (PNL).
The most general class with known sufficient conditions for identifiability of the graph is the Post-Nonlinear (PNL) model (Zhang & Hyvärinen, 2009), in which
$X_{i}\coloneqq g_{i}\!\big(f_{i}(\operatorname{PA}_{i})+S_{i}\big),\quad i=1,\ldots,d,$ (31)
with $f_{i}$ and $g_{i}$ both potentially nonlinear, $g_{i}$ invertible, and mutually independent noises.
Location Scale Noise Model (LSNM)
The LSNM (Immer et al., 2022) extends ANMs by allowing heteroscedastic noise as follows:
$X_{i}\coloneqq f_{i}(\operatorname{PA}_{i})+g_{i}(\operatorname{PA}_{i})\,S_{i},\quad i=1,\ldots,d,$ (32)
where $f_{i}$ and $g_{i}>0$ may be nonlinear and noise terms are jointly independent with zero mean and unit variance.
C.3 Detailed pseudocode of Algorithm 1
Algorithm˜2 provides a detailed pseudocode of the algorithm adopted in our experiments of Section˜5, and sketched in Algorithm˜1.
Data: $\mathcal{D}\in\mathbb{R}^{k\times n\times d}$
// $\forall$ env: $n$ d-dimensional observations.
    $I_{1},I_{2}\subset[k]$
// Set of indices splitting the environments in two groups
Result: Estimate of $\operatorname{supp}J_{\mathbf{f}^{-1}}$
$\widehat{S}\leftarrow\operatorname{score\_estimate}(\mathcal{D})\in\mathbb{R}^{k\times n\times d}$
$\widehat{H}\leftarrow\operatorname{hess\_estimate}(\mathcal{D})\in\mathbb{R}^{k\times n\times d\times d}$
$\operatorname{mean\_pairs\_idxs}\in\mathbb{R}^{e\times 2}$
// Pair of indices corresponding to observations at the mean
// For each environment $e$, find $i$ s.t. $\mathbf{f}^{-1}(X[e,i])\approx\mu_{\mathbf{S}}$
for e=1,,ke=1,...,k do
 $\Delta_{X}\in\mathbb{R}^{n\times n}$
 // norm of the difference of observations from distinct envs
 $\operatorname{pairs}\in\mathbb{N}^{n}$
 // Pair of indices $i,j$ such that $X[0,i]\approx X[e,j]$
 $\operatorname{score\_diffs}\leftarrow+\infty\in\mathbb{R}^{n}$
 // Container for norm of the differences in the score
 for i = 1,…,n do
    for j=1,…,n do
[…]
      end for
    $j\leftarrow\operatorname*{arg\,min}\Delta_{X}[i]$
    $\operatorname{pairs}[i]\leftarrow j$
    // $X[0,i]\approx X[e,j]$
[…]
   end for
 $m\leftarrow\operatorname*{arg\,min}\operatorname{score\_diffs}$
 // Paired observations between envs $(0,e)$ s.t. score diff. $\approx 0$.
 $\operatorname{mean\_pairs\_idxs}[e]\leftarrow m,\operatorname{pairs}[m]$
 // The score diff. vanishes when source == mean
end for
// Difference of Hessians at the mean (i.e. Equations˜9 and 10)
$\widehat{H}_{\textnormal{diffs}}\leftarrow 0\in\mathbb{R}^{2\times d\times d}$
for $\ell=1,2$ do
 for $e\in I_{\ell}$ do
    $m_{1},m_{e}\leftarrow\operatorname{mean\_pairs\_idxs}[e]$
    $\Delta_{H}=\widehat{H}[0,m_{1}]-\widehat{H}[e,m_{e}]$
    $\widehat{H}_{\textnormal{diffs}}[\ell]\leftarrow\widehat{H}_{\textnormal{diffs}}[\ell]+\Delta_{H}$.
   end for
end for
$M\leftarrow\widehat{H}_{\textnormal{diffs}}^{-1}[1]\widehat{H}_{\textnormal{diffs}}[2]\approx J_{\mathbf{f}}\Omega_{1}^{-1}\Omega_{2}J_{\mathbf{f}^{-1}}$
// $H_{\textnormal{diffs}}[\ell]\approx J_{\mathbf{f}^{-1}}^{T}\Omega_{\ell}J_{\mathbf{f}^{-1}}$, by Equations˜9 and 10
$\widehat{J}_{\mathbf{f}^{-1}}\leftarrow\textnormal{diagonalize}(M)\approx J_{\mathbf{f}^{-1}}DP$
return $\operatorname{supp}\left(\widehat{J}_{\mathbf{f}^{-1}}P^{-1}\right)$
// $P$ can be found using the acyclicity of the causal graph.
Algorithm 2 Estimating $\operatorname{supp}J_{\mathbf{f}^{-1}}$ from the data
C.4 Experiments beyond Gaussianity
In this section, we present additional experimental results on bivariate graphs underlying synthetically generated structural causal models. The causal mechanisms are the same already described in Section˜5.1. The difference, here, is that we generate the independent sources from a Gamma distribution, which violates the assumptions of our theory. We sample the scale parameter $\theta\sim U(1.75,2.25)$, and consider two different parameterizations of the shape $\alpha$ of the base environments: in the first case, $\alpha\sim U(0.5,1)$; in the second case $\alpha\sim U(2,2.5)$. What makes the Gamma density interesting it that it can be flexibly modified by changing the values of its parameters, as shown in Figure˜2.
Gamma distribution with no vanishing gradient.
Figure˜2(a) illustrate how the Gamma density function varies at $\alpha=1$ and different values of $\theta$. It is interesting to note that the gradient of the density function never vanishes, making this setup adversarial to the assumptions of Theorem˜1. In line with this, in Figure˜3 we see that generally our algorithm struggles to infer the causal direction for this class of structural causal models.
Gamma distribution with vanishing gradient.
Figure˜2(b) illustrates how the Gamma density function varies at $\alpha=2$ and different values of $\theta$. We can see that, in this case, the density achieves a maximum: we point to our analysis in Section˜4.1 (the paragraph Theorem˜1 beyond Gaussianity), where we discuss when and why it is reasonable to expect that Theorem˜1 extends to any source distribution that achieves a maximum or minimum in the interior of its domain. A word of caution is needed: despite the fact that the Gamma density with $\alpha\in[2,2.5]$ does have a vanishing gradient, the points of the domain at which the critical values occur are not preserved by our rescaling interventions (as is clear by inspection of Figure˜2(b)). Hence, the requirements of the Theorem˜1 are not fully met (where it’s implicit that the rescaling interventions do not change the location of the modes): this makes the experiments of Figure˜4 an interesting challenge for our algorithm. The outcomes are exciting: we see that increasing the number of available environments, despite the assumption violations, imposes enough constraints to infer the causal direction in the majority of the experimental setups with $\approx 80\%$ accuracy. This is of double interest: first, we have some empirical evidence supporting the hypothesis that our theory can be extended beyond Gaussianity. Second, we see that this seems to be achieved thanks to the constraints from many environments, in contrast with what we observe when experiments are run on SCMs with Gaussian noise (Figure˜1), where increasing environments do not translate into better accuracy. These empirical findings, despite being preliminary, should provide an incentive to pursue identifiability theory beyond Gaussianity.
(a) Gamma density with $\alpha=1$.
(b) Gamma density with $\alpha=2$.
Figure 2: We plot the Gamma density for different values of shape and scale. The left plot fixes the shape $\alpha=1$; the right plot fixes $\alpha=2$. We let $\theta$ vary to illustrate how the distribution changes between the rescaling environments of our experiments. We note that for $\alpha=1$ the density doesn’t have a finite critical point.
Figure 3: Average SHD (0 is best, 11 is worst) achieved by Algorithm˜1 over 5050 seeds on binary graphs. The sources are sampled from a gamma distribution with $\alpha\in[0.5,1]$. In line with our theory, when the sources are generated according to a density that doesn’t have critical points, our algorithm generally fails to infer the causal direction.
Figure 4: Average SHD (0 is best, 11 is worst) achieved by Algorithm˜1 over 5050 seeds on binary graphs. The sources are sampled from a gamma distribution with $\alpha\in[2,2.5]$, which guarantees at least one point where the gradient of the log-likelihood vanishes (see Figure˜2(b)). Interestingly, this appears to enable accurate inference of the causal graph when the number of environments increases.
Appendix D Additional content
In this section, we collect some useful results and notes relevant to the main paper.
D.1 Graph theory
Directed graphs and DAGs.
Let $X_{1},\ldots,X_{d}$ be a vector of random variables. A graph $\mathcal{G}=(\{X_{i}\}_{i}^{d},E)$ consists of a vertex set $\{X_{i}\}_{i}^{d}$ and an edge set $E$. We recall a few basic notions for directed graphs.
A directed edge $X_{i}\to X_{j}$ indicates that $X_{i}$ is a parent of $X_{j}$ (and $X_{j}$ a child of $X_{i}$). $\operatorname{PA}_{i}\subset[d]$ denotes the index of the parent nodes of $X_{i}$ in the graph $\mathcal{G}$, $\operatorname{CH}_{i}\subset[d]$ denotes the children. A path in $\mathcal{G}$ is a sequence of at least two distinct vertices $\pi=X_{i_{1}},\ldots,X_{i_{m}}$ such that each consecutive pair $X_{i_{k}}$ and $X_{i_{k+1}}$ is joined by an edge for $k=1,\ldots,m-1$. If every edge along the path is oriented forward, $X_{i_{k}}\to X_{i_{k+1}}$, we call it a directed path; then $X_{i_{1}}$ is an ancestor of $X_{i_{m}}$ and $X_{i_{m}}$ a descendant of $X_{i_{1}}$.
D.2 From SCM to ICA models
Equation˜2 claims that structural causal models can be expressed in the form of ICA models. Here, we show how this can be achieved. Consider a set of causal variables $\mathbf{X}=(X_{i})_{i=1}^{d}$, and without loss of generality, assume that the causal order is 1,,d1,...,d. According to Equation˜1, for each i=1,,d,i=1,...,d, we have:
$X_{i}:=F_{i}(\mathbf{X}_{\operatorname{PA}_{i}},S_{i}),$
with $\mathbf{S}=(S_{i})_{i=1}^{d}$ the vector of mutually independent noise terms. An inductive argument shows the existence of a function $f_{i}:\mathbf{S}_{\operatorname{AN}_{i}}\mapsto X_{i}$, where $\operatorname{AN}_{i}$ denotes the indices of the ancestor nodes of $X_{i}$ in the causal graph. Given the causal order 1,,d1,...,d, the base case is given for $X_{1}:=F_{1}(S_{1})$, such that $f_{1}:=F_{1}$. The inductive step is as follows: assume that there is $n<d$ such that $X_{i}=f_{i}(\mathbf{S}_{\operatorname{AN}_{i}},S_{i})$ for all i=1,,ni=1,...,n. Then, there is a map $\mathbf{S}_{[n]}\mapsto\mathbf{X}_{[n]}$. The causal order 1,,d1,...,d implies $\operatorname{AN}_{n+1}\subset[n]$, so that there is a map $\mathbf{S}_{[n]}\mapsto\mathbf{X}_{\operatorname{AN}_{n+1}}$: given that $\operatorname{PA}_{n+1}\subseteq\operatorname{AN}_{n+1}$, there is a map $g:\mathbf{S}_{[n]}\mapsto\mathbf{X}_{\operatorname{PA}_{n+1}}$: from the structural equation $X_{n+1}:=F_{n+1}(\mathbf{X}_{\operatorname{PA}_{n+1}},S_{n+1})=F_{n+1}(g(\mathbf{S}_{\operatorname{AN}_{n+1}}),S_{n+1})$, we conclude that there is $f_{n+1}:\mathbf{S}_{\operatorname{AN}_{n+1}},S_{n+1}\mapsto\mathbf{X}_{n+1}$. Then, we define $\mathbf{f}:=(f_{i})_{i=1}^{d}$ and find
$\mathbf{X}=\mathbf{f}(\mathbf{S}).$
An important note is that the DAG structure of the causal graph is reflected in the Jacobian of the mixing function $\mathbf{f}$, which can be shown to be lower triangular.
D.3 Hessian of the log-density of independent random variables
In the main paper we mention that the $\Omega_{1},\Omega_{2}$ matrices defined in Equation˜8 are diagonal; here, we discuss why this is true. More generally, it is well known that for a vector of independent random variables $\mathbf{Z}\in\mathbb{R}^{d}$ with density $p$, the following holds:
[…]
D.4 Measure theoretic arguments in support of the assumptions
First, we show that ˜5 generically holds.
Proposition 3 (˜5 holds almost surely).
Let $L_{i}=\mathrm{diag}(\lambda_{1}^{i},\dots,\lambda_{d}^{i})$ and $\lambda_{j}^{i}\neq 0$, i=1,,ki=1,...,k. Assume the joint law of the array $\Lambda=(\lambda_{j}^{i})_{j\in[d],\,i\in[k]}$ is absolutely continuous with respect to Lebesgue measure on $\bigl(\mathbb{R}\setminus\{0\}\bigr)^{dk}$. Then, with probability one over the draw of $\Lambda$: for every $j\in[d]$,
$\sum_{i\in[k]}\frac{1}{(\lambda_{j}^{i})^{2}}\ \neq\ k.$
Proof.
Fix $j\in[d]$. Write $k=|I_{l}|$ and $\lambda:=(\lambda_{j}^{i})_{i\in[k]}\in(\mathbb{R}\setminus\{0\})^{k}$. Consider the smooth map $F:\,(\mathbb{R}\setminus\{0\})^{k}\to\mathbb{R}$,
$F(\lambda)\ =\ \sum_{r=1}^{k}\lambda_{r}^{-2}\ -\ k.$
Its gradient is $\nabla F(\lambda)=(-2\lambda_{1}^{-3},\dots,-2\lambda_{k}^{-3})\neq 0$ on the domain, so 0 is a regular value. By the regular level–set theorem, $F^{-1}(0)$ is a (k1)(k-1)-dimensional embedded submanifold of $\mathbb{R}^{k}$ and hence has Lebesgue measure zero. Because the $k$-tuple $\lambda=(\lambda_{j}^{i})_{i\in[k]}$ has a distribution absolutely continuous with respect to Lebesgue measure, we get
$\mathbb{P}\!\left(\sum_{i\in I_{l}}\frac{1}{(\lambda_{j}^{i})^{2}}=k\right)=0.$
Taking the finite union over $j=1,\dots,d$ preserves measure zero, so with probability one none of these equalities occurs. ∎
Next, we show that the assumption of pairwise distinct $\{(\Omega_{1}\Omega_{2}^{-1})_{i}i\}_{i\in[d]}$ elements (definition at Equation˜8) generically holds.
Proposition 4 (Pairwise distinct diagonal ratios hold almost surely).
Let $I_{1},I_{2}\subset[k]\geq 3$. For each environment $i$ let $L_{i}=\mathrm{diag}(\lambda^{i}_{1},\dots,\lambda^{i}_{d})$ with $\lambda^{i}_{j}\neq 0$. Assume the joint law of the array $\Lambda=(\lambda_{j}^{i})_{j\in[d],\,i\in[k]}$ is absolutely continuous with respect to Lebesgue measure on $\bigl(\mathbb{R}\setminus\{0\}\bigr)^{dk}$. Suppose moreover that $\Omega_{\ell}$ is diagonal with entries
$(\Omega_{\ell})_{jj}\;=\;\frac{1}{\sigma_{j}^{2}}\Bigl(\sum_{i\in I_{\ell}}(\lambda^{i}_{j})^{-2}-|I_{\ell}|\Bigr)\neq 0.\qquad\ell\in\{1,2\},\ j\in[d],$
Then, with probability one over the draw of $\Lambda$, $\Omega_{1}$ is invertible and the diagonal entries of $\Omega_{1}^{-1}\Omega_{2}$ are pairwise distinct.
Proof.
Write
$(\Omega_{1}^{-1}\Omega_{2})_{jj}\;=\;\frac{\sum_{i\in I_{2}}(\lambda^{i}_{j})^{-2}-|I_{2}|}{\sum_{i\in I_{1}}(\lambda^{i}_{j})^{-2}-|I_{1}|}\;=:\;\frac{B_{j}}{A_{j}},\qquad A_{j}:=\sum_{i\in I_{1}}(\lambda^{i}_{j})^{-2}-|I_{1}|,\ \ B_{j}:=\sum_{i\in I_{2}}(\lambda^{i}_{j})^{-2}-|I_{2}|.$
By Proposition˜3, $A_{j}\neq 0$ and $B_{j}\neq 0$ for all $j$ with probability one, such that $\Omega_{1}$ is invertible.
Fix $j\neq\ell$. The collision event $(\Omega_{1}^{-1}\Omega_{2})_{jj}=(\Omega_{1}^{-1}\Omega_{2})_{\ell\ell}$ is equivalent to
$\frac{B_{j}}{A_{j}}=\frac{B_{\ell}}{A_{\ell}}\quad\Longleftrightarrow\quad F_{j\ell}(\Lambda):=A_{j}B_{\ell}-A_{\ell}B_{j}=0.$
Let $t^{i}_{h}:=(\lambda^{i}_{h})^{-2}$ and view $F_{j\ell}$ as a smooth function of the $2k$ variables $\{t^{i}_{j}\}_{i\in[k]}\cup\{t^{i}_{\ell}\}_{i\in[k]}$. For any fixed $i_{0}\in I_{1}$,
$\frac{\partial F_{j\ell}}{\partial t^{\,i_{0}}_{j}}=\frac{\partial A_{j}}{\partial t^{\,i_{0}}_{j}}\,B_{\ell}-A_{\ell}\,\frac{\partial B_{j}}{\partial t^{\,i_{0}}_{j}}=1\cdot B_{\ell}-A_{\ell}\cdot 0=B_{\ell}.$
Since $B_{\ell}\neq 0$, we have $\nabla F_{j\ell}\neq 0$ on the set under consideration, so 0 is a regular value of $F_{j\ell}$. By the regular level-set theorem, the set $\{F_{j\ell}=0\}$ is a (2j1)(2j-1)-dimensional embedded submanifold of $\mathbb{R}^{2k}$, hence it has Lebesgue measure zero. Because the law of $\Lambda$ is absolutely continuous w.r.t. the Lebesgue measure,
$\mathbb{P}\left((\Omega_{1}^{-1}\Omega_{2})_{jj}=(\Omega_{1}^{-1}\Omega_{2})_{\ell\ell}\right)=0.$
Taking the finite union over all pairs $j\neq\ell$ yields that, with probability one, no two diagonal entries coincide; that is, $\{(\Omega_{1}^{-1}\Omega_{2})_{jj}\}_{j=1}^{d}$ are pairwise distinct. ∎